OpenStudy (anonymous):

What is the x-coordinate of the x-intercept of the line that passes through points (-5, 3) and (1, -3)?

OpenStudy (lgbasallote):

first step: find the equation of that line. do you know how?

OpenStudy (anonymous):

no

OpenStudy (lgbasallote):

use the formula \[\Large y -y_1 = \frac{y_2 - y_1}{x_2 - x_1} (x-x_1)\] do you know what those notations mean?

OpenStudy (anonymous):

yes

OpenStudy (lgbasallote):

wonderful

OpenStudy (lgbasallote):

tell me when you get a result

OpenStudy (anonymous):

its y2-y1 first right?

OpenStudy (lgbasallote):

what do you mean?

OpenStudy (anonymous):

is the answer -1?

OpenStudy (lgbasallote):

how did you get that?

OpenStudy (anonymous):

-3-3=-3-3/1- -5(1- -5)

OpenStudy (lgbasallote):

noo...

OpenStudy (anonymous):

where'd i go wrong

OpenStudy (lgbasallote):

it's y - y1 and x - x1 NOT y2 - y1 and x2 -x1

OpenStudy (lgbasallote):

look at the formula i wrote

OpenStudy (anonymous):

well whats y and x?

OpenStudy (lgbasallote):

nothing..you just put them as y and x

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

y-3=-6/6(x+5)

OpenStudy (lgbasallote):

correct simplify it

OpenStudy (anonymous):

y-3=x+5

OpenStudy (anonymous):

*-x+5

OpenStudy (lgbasallote):

good y - 3 = -x - 5 now simplify it more

OpenStudy (anonymous):

how?

OpenStudy (lgbasallote):

add 3 to both sides

OpenStudy (anonymous):

y=-x-2

OpenStudy (lgbasallote):

good. now let y = 0 and solve for x

OpenStudy (anonymous):

x=2

OpenStudy (lgbasallote):

that is now the x coordinate of the x intercept of the line

OpenStudy (anonymous):

thanks

OpenStudy (lgbasallote):

welcome